# MuJo: Multimodal Joint Feature Space Learning for Human Activity Recognition

### v1

@@ Abstract
Human Activity Recognition is a longstanding problem in AI with applications in a broad range of areas: from healthcare, sports and fitness, security, and human computer interaction to robotics. The performance of HAR in real-world settings is strongly dependent on the type and quality of the input signal that can be acquired. Given an unobstructed, high-quality camera view of a scene, computer vision systems, in particular in conjunction with foundational models (e.g., CLIP), can today fairly reliably distinguish complex activities. On the other hand, recognition using modalities such as wearable sensors (which are often more broadly available, e.g, in mobile phones and smartwatches) is a more difficult problem, as the signals often contain less information and labeled training data is more difficult to acquire. In this work, we show how we can improve HAR performance across different modalities using multimodal contrastive pretraining. Our approach MuJo (Multimodal Joint Feature Space Learning), learns a multimodal joint feature space with video, language, pose, and IMU sensor data. The proposed approach combines contrastive and multitask learning methods and analyzes different multitasking strategies for learning a compact shared representation. A large dataset with parallel video, language, pose, and sensor data points is also introduced to support the research, along with an analysis of the robustness of the multimodal joint space for modal-incomplete and low-resource data. On the MM-Fit dataset, our model achieves an impressive Macro -Score of up to 0.992 with only 2% of the train data and 0.999 when using all available training data for classification tasks. Moreover, in the scenario where the MM-Fit dataset is unseen, we demonstrate a generalization performance of up to 0.638.
@@ 1 Introduction
Perceiving and interpreting human activity is a core functionality of intelligent systems with applications in a broad range of areas: from healthcare Tan et al. (2021) through sports and fitness Host and Ivašić-Kos (2022); Nadeem et al. (2020), security Sunil et al. (2021) to robotics Piyathilaka and Kodagoda (2015). A basic capability needed to facilitate such functionality is the identification of concrete human actions from the system’s sensory input, referred to as Human Activity Recognition (HAR) Hamad et al. (2023). In general, HAR is a difficult problem, especially when applied to uncontrolled real-life environments, where it must contend with the variability of human actions and the open-ended variations in environmental conditions. In such settings, the performance of HAR systems is strongly dependent on the available input modalities. Recent advances in Computer Vision, particularly with Large Vision Language Models (VLMs) Radford et al. (2021); Achiam et al. (2023), have enabled impressive performance when given sufficiently high-quality images. In many cases, merely feeding images to publicly available tools such as GPT-4 can result in accurate description of scenes involving complex and subtle activities. However, in real-life settings, such images are not always available; either due to a lack of cameras or because of privacy concerns. Much more broadly available input modalities are wearable sensors, in particular, Inertial Measurement Units (IMUs), which are ubiquitous in smartphones, smartwatches or even earbuds, all of which people carry with them on a daily basis. Unfortunately, achieving acceptable recognition performance with such sensors is a much more difficult problem due to two factors:
@@
Information ambiguity: An IMU (which is essentially a motion sensor) in a smartwatch on the wrist or in a smartphone in a pocket contains much less detailed information about the user’s activity than a high-quality picture.
2.
Lack of training data: The success of VLMs is built on the online availability of massive amounts of text/image data pairs, which can be used for training aligned representations. Although large amounts of unlabeled IMU data are increasingly becoming available Chan et al. (2024), labeled training data remains scarce.
The work presented in this paper, which aims to address the issues outlined previously, is founded on three key observations:
1.
While the input modalities available during inference—the time when the system is utilized in the real world—may be constrained, the input modalities during training can be selected with much greater flexibility.
@@
Significant progress has recently been made in generating synthetic data for a variety of sensors from videos. For example, IMUTube Kwon et al. (2021) generates virtual gyroscope and accelerometer data from videos, while pose extraction methods like OpenPose Cao et al. (2017) and VideoPose3D Pavllo et al. (2019) are used to generate skeletal data of human activities. This enables the creation of synchronised video/sensor training data and through the alignment of video with text, it facilitates the synchronization of sensor data with textual descriptions.
Based on the above considerations, we propose a multimodal contrastive pretraining method that creates a joint representation between videos, video-derived poses, synthetic IMU data created from videos using IMUTube Kwon et al. (2021), and textual descriptions. Our approach, which we refer to as MuJo, aims to utilize the complementary information of each modality to improve the performance for both unimodal and multimodal inference. In particular, we show that our method dramatically reduces the need for labeled training data for downstream tasks. Thus, on the well-known MM-Fit dataset Strömbäck et al. (2020), when using all available training data, our method improves the recognition rate (Macro -Score) from 0.973 to 0.999 in the multimodal case (encompassing all of the above modalities), from 0.957 to 0.988 for poses, and from 0.727 to 0.863 for video alone. For sensor data, the performance of our method is slightly worse (0.798 as opposed to 0.822), which is due to the fact that MM-Fit uses real, rather than simulated sensor data. However, when using just 2% of the available data for training, the baseline with sensors has an -Score of just 0.378 whereas our systems achieves a reasonable 0.627 (see section 6.2 for details). The multimodal -Score is 0.819 in the baseline, while our method raises it to 0.992.
@@
Development of a large fitness dataset111Available after publication at link. with parallel video, language (i.e., labels of video clips, captions uttered by the instructor in the video clip, and precise descriptions of the exercises generated by GPT-3.5), pose, and simulated IMU data points.
Learning of an effective joint space for the modalities of text, video, pose, and sensor by leveraging the contrastive loss.
Analysis of usefulness and generalizability of multimodal joint space for each modality and their combination, as well as the benefits of pretraining when a low amount of data is available for the target task.
@@ 2 Related Work
2.1 Human Activity Recognition
Unimodal activity recognition.
Unimodal activity recognition has been extensively studied in HAR. Various approaches have been proposed to address this task, and they can be broadly categorized into two groups: hand-crafted feature-based and deep learning-based methods. Hand-crafted feature-based methods rely on extracting low-level features from the input data and designing a classifier to recognize activities. For instance, in accelerometer-based activity recognition Kwapisz et al. (2011), features such as mean, standard deviation, and energy of the acceleration signal are often used to capture the characteristics of different activities. Popular classifiers used in this approach include decision trees, support vector machines, and k-nearest neighbors. Deep learning-based methods, on the other hand, use neural networks to learn features and recognize activities from raw input data. Convolutional neural networks (CNNs) and recurrent neural networks (RNNs) are two popular types of neural networks used in this approach. CNNs are commonly used for image-based activity recognition Strömbäck et al. (2020); Ordóñez and Roggen (2016), being combined with RNNs in some cases Ordóñez and Roggen (2016). Despite the success of unimodal activity recognition, it is often limited by the fact that a single modality of input data may only capture some of the necessary information about the activity, leading to lower recognition accuracy. Therefore, multimodal activity recognition, which combines information from multiple sources, has gained increasing attention in recent years.
Multimodal activity recognition.
Multimodal activity recognition recognizes human activities using information from multiple modalities, such as IMU data, audio, and video Mekruksavanich and Jitpattanakul (2022); Ijaz et al. (2022). These features can then be combined using fusion methods such as early fusion or late fusion Duhme et al. (2022). Some studies have also focused on domain adaptation, where the goal is to recognize activities in a new environment using data from a different environment Plananamente et al. (2022). In such cases, transfer learning techniques can adapt the model to the new environment using a small amount of labeled data. Overall, multimodal activity recognition has shown promising results, with improved accuracy compared to unimodal approaches. However, challenges in data fusion, feature extraction, and model selection still need to be addressed to improve the performance of multimodal activity recognition systems Islam et al. (2022).
2.2 Representation Learning for Human Activities
@@
Different modalities, e.g., natural language, visual inputs, or sensor signals, often complement a common concept. A key challenge of multimodal feature representation is exploring efficient methods for multimodal fusion with the given modalities to avoid missing modal features. We can differentiate the representation learning methods in two aspects: learning objectives and fusing the learned features in a joint space. In learning objectives, various approaches have been proposed to learn multimodal representations by formulating proxy tasks such as reconstruction, completion Seo et al. (2021); Sun et al. (2019), matching/alignment Li et al. (2020); Miech et al. (2020), and ordering tasks Li et al. (2020). Generating one feature vector with multimodal inputs for a common concept is a crucial problem of multimodal representation learning. Simple methods like concatenating Hu et al. (2018); Strömbäck et al. (2020) learned uni-modal feature vectors have been extensively explored to enhance the performance of downstream tasks. These methods involve combining feature vectors from a single modality, such as text or image, to create more informative representations that can be used for various tasks. Thereafter, Zadeh et al. proposed a tensor fusion network in Zadeh et al. (2017). Lu et al. proposed a co-attention mechanism based on the transformer architecture in Lu et al. (2019) for efficient cross-modality learning.
@@
Contrastive Learning is as an efficient self-supervised framework applied across multiple domains, which learns similar/dissimilar representations from data that are organized into similar/dissimilar pairs. Recently, Radford et al. presented a novel approach, known as CLIP Radford et al. (2021), to jointly represent images and text descriptions by leveraging contrastive loss with the similarity of image-text pairs. CLIP learns two feature spaces with two encoders, i.e., image and text encoders, then projects the learned features into a shared latent space with contrastive loss. Contrastive learning, a joint feature space of text-image pairs, has been extended to other modalities. For example, video-text data has attracted studies on shared space of video-text features Xu et al. (2021); Xue et al. (2022), text-to-video generation Singer et al. (2022). Thereafter, MotionCLIP Tevet et al. (2022) trained an encoder to find the proper embedding of an input sequence in CLIP space, and a decoder that generates the most fitting motion to a given CLIP space. Following that, IMU2CLIP Moon et al. (2022) has been proposed to learn a joint space to align Inertial Measurement Unit (IMU) motion sensor recordings with video and text, by projecting them into the joint representation space.
@@
Studies in action recognition Aggarwal and Ryoo (2011); Li et al. (2020); Wang et al. (2021); Sun et al. (2022); Morshed et al. (2023) can also significantly enhance the field of recognition in fitness activities, despite action recognition emphasis on a broad spectrum of actions such as “cutting the potato” and “cleaning the drawer”, among others. Certain investigations in action recognition also endeavor to acquire knowledge of the feature space through contrastive learning methodologies, exemplified by initiatives such as HERO Li et al. (2020) and ActionCLIP Wang et al. (2021).
@@
The achievements discussed in the aforementioned studies have shown contrastive learning has a high impact on extracting a joint space with similar and dissimilar pairs. Hence, the contrastive learning methodology employed in this study is utilized to acquire knowledge of the feature space pertaining to fitness activities. Contrary to previous work, we extend this line of work on contrastive learning to a unique multimodal setting that utilizes more than two modalities, such as video, IMU sensors, poses, and text, within a multitask training manner in an end-to-end training paradigm similar to ImageBind Girdhar et al. (2023). Thus, our model and data provide a cross-modal classification opportunity.
3 Dataset
3.1 Video Collection
@@
YouTube provides chapters within a video, each with its preview, providing information and context for different parts of the video. Video owner can manually add chapters to their uploaded videos or rely on automatic chapter generation. We manually collect 1212 fitness videos with pre-annotated video chapters. Additionally, we retrieve the captions automatically generated by YouTube. Our videos include only one tutor who teaches how to do fitness exercises (e.g., “squat with dumbbells”, “sit-ups”, “side deep squats”, etc.). Thus, the exercise videos provide instructions for the activities while showing the activities clearly. The annotation does not include predefined categories or labels. Each video owner annotates the activity segment of the videos with their personal preferences, such as “deep squats with dumbbells”, “high jump squats”, etc., instead of only “squats”. Given the temporal boundaries of the activity within each video, we split the videos in shorter activity clips along with their corresponding chapters and caption segments.
Since workout exercises are typically short, activity videos that are too long often contain multiple exercises and, thus, incorrect chapter labels. Hence, we only keep exercises shorter than 2 minutes and discard more extended exercises. After that, the dataset consists of 10,695 (video, chapter, caption) samples.
3.2 Dataset Construction
@@
We convert all videos to 720p resolution and 50 frames per second via FFmpegs adaptive overlapped block motion compensation with bidirectional motion estimation. Then, we extract simulated sensor data, poses, as well as a variety of video and text features. Video features are extracted with HERO Li et al. (2020) utilizing CLIP Radford et al. (2021) and S3D Xie et al. (2018) embeddings. On the textual side, we apply several normalization strategies to the raw chapters such as replacing abbreviations within the text (e.g., changing “l” to “left” and “r” to “right”) as well as removing stop words, numbers, and a few special symbols from a predefined list. Further, we use GPT 3.5 to generate detailed descriptions of the exercises. We then extract multiple text features with different models, including OpenAI’s text-embedding-ada-002 embedding model (ADA), CLIP, BERT Devlin et al. (2018), as well as language features and motion tokens extracted with Text2Motion Transformer Guo et al. (2022). We apply IMUTube Kwon et al. (2021) to generate virtual gyroscope and accelerometer data for sensors placed on all 16 joint nodes. We also use the implementation of IMUTube to extract 3D poses from the videos, which utilizes OpenPose followed by VideoPose3D for 3d pose lifting.
Since IMUTube fails to extract sensor data for 884 videos, we exclude these sequences. We use the remaining 9811 exercises as our dataset and split it into train (8,518 videos), validation (946 videos) and test (347 videos) sets. Since workout session videos typically start with an introductory sequence and conclude with a farewell, we discard the initial and final 15%222This value has proven to be effective in experimental investigations. of each video. Despite these filtering steps, we receive a large dataset totaling around 78 hours in length. Instances are generated using sliding windows of 100 frames in length (i.e., 2 seconds) with 50 frames overlap from the videos. We found this size suitable for fitness exercises as a 2 second window generally captures enough meaningful information regarding the subject’s movements.
In summary, our dataset comprises a large collection of short fitness activity clips containing video features, poses, synthetic IMU data, chapter labels, captions, and GPT-generated descriptions as well as several types of features extracted from these text modalities.

### v2

@@ Abstract
Human Activity Recognition (HAR) is a longstanding problem in AI with applications in a broad range of areas, including healthcare, sports and fitness, security, and more. The performance of HAR in real-world settings is strongly dependent on the type and quality of the input signal that can be acquired. Given an unobstructed, high-quality camera view of a scene, computer vision systems, in particular in conjunction with foundation models, can today fairly reliably distinguish complex activities. On the other hand, recognition using modalities such as wearable sensors (which are often more broadly available, e.g., in mobile phones and smartwatches) is a more difficult problem, as the signals often contain less information and labeled training data is more difficult to acquire. To alleviate the need for labeled data, we introduce our comprehensive Fitness Multimodal Activity Dataset (FiMAD) in this work, which can be used with the proposed pre-training method MuJo (Multimodal Joint Feature Space Learning) to enhance HAR performance across various modalities. FiMAD was created using YouTube fitness videos and contains parallel video, language, pose, and simulated IMU sensor data. MuJo utilizes this dataset to learn a joint feature space for these modalities. We show that classifiers pre-trained on FiMAD can increase the performance on real HAR datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH. For instance, on MM-Fit, we achieve an Macro -Score of up to 0.855 when fine-tuning on only 2% of the training data and 0.942 when utilizing the full training set for classification tasks. We have compared our approach to other self-supervised ones and showed that, unlike them, ours can consistently improve on the baseline network performance as well as provide a better data-efficiency.
@@ 1 Introduction
Perceiving and interpreting human activities is a core functionality of intelligent systems with applications in a broad range of areas, including healthcare Tan et al. (2021), sports and fitness Host and Ivašić-Kos (2022); Nadeem et al. (2020), and security Sunil et al. (2021), among others. A basic capability needed to facilitate such functionality is the identification of concrete human actions from the system’s sensory input, referred to as Human Activity Recognition (HAR). In general, HAR is a difficult problem, especially when applied to uncontrolled real-life environments, where it must contend with the variability of human actions and the open-ended variations in environmental conditions. In such settings, the performance of HAR systems is strongly dependent on the available input modalities. Recent advances in Computer Vision, particularly with Large Vision-Language Models (LVLMs) Radford et al. (2021); Achiam et al. (2023), have enabled impressive performance when given sufficiently high-quality images. In many cases, merely feeding images to publicly available tools such as GPT-4 can result in accurate description of scenes involving complex and subtle activities. However, in real-life settings, such images are not always available; either due to a lack of cameras or because of privacy concerns. Much more broadly available input modalities are wearable sensors, in particular, Inertial Measurement Units (IMUs), which are ubiquitous in smartphones, smartwatches or even earbuds, all of which people carry with them on a daily basis. Unfortunately, achieving acceptable recognition performance with such sensors is a much more challenging problem due to two factors, the first one being information ambiguity. An IMU (a motion sensor) in a smartwatch on the wrist or in a smartphone in a pocket contains much less detailed information about the user’s activity than a high-quality picture. This problem is heightened by the second factor, lack of training data. The success of LVLMs is built on the online availability of massive amounts of text/image data pairs, which can be used for training aligned representations. Although large amounts of unlabeled IMU data are increasingly becoming available Chan et al. (2024), labeled training data remains scarce.
The work presented in this paper, which aims to address these issues, is founded on three key observations:
@@
While the input modalities availability during inference—the time when the system is utilized in the real world—may be constrained, the input modalities during training can be selected with much greater flexibility.
@@
Significant progress has recently been made in generating synthetic data for a variety of sensors from videos. For example, IMUTube Kwon et al. (2021) generates virtual gyroscope and accelerometer data from videos, while pose extraction methods like OpenPose Cao et al. (2017) and VideoPose3D Pavllo et al. (2019) are used to generate skeletal data of human activities. This enables the creation of synchronized video/sensor training data and through the alignment of video with text, it facilitates the synchronization of sensor data with textual descriptions.
Based on the above considerations, we propose a Multimodal Joint Feature Space Learning method (MuJo) that learns a joint representation between videos, video-derived poses, virtual IMU data generated from videos, and textual descriptions by employing pair-wise contrastive learning, as depicted in Figure 1. MuJo aims to utilize the complementary information of each modality to improve the performance for both unimodal and multimodal inference. In particular, we show that our method dramatically reduces the need for labeled training data for downstream tasks. For that, we provide extensive evaluation on four publicly available HAR datasets, namely MM-Fit Strömbäck et al. (2020), MyoGym Koskimäki et al. (2017), MotionSense Malekzadeh et al. (2018), and MHEALTH Banos et al. (2015). For instance, on the MM-Fit dataset in the multimodal case (encompassing all of the aforementioned modalities as input), our method improves the recognition rate (Macro -Score) from 0.897 to 0.942 when using all available training data. Even when utilizing only 2% of the training data, it enhances the recognition rate from 0.790 to 0.855.
@@
1.
Development of the Fitness Multimodal Activity Dataset (FiMAD)111Available after publication at link., a comprehensive dataset containing parallel video, language (including video clip labels, instructor utterances, and detailed exercise descriptions generated by GPT-3.5 Turbo), pose data, and simulated IMU data points.
2.
Pe-training of MuJo, a multimodal joint space model, using the FiMAD dataset. MuJo integrates four different modalities: text, video, pose, and virtual sensor data, leveraging contrastive learning inspired by CLIP Radford et al. (2021).
3.
Evaluation of the utility and generalizability of MuJo for real-world accelerometer data on the MM-Fit, MyoGym, MotionSense, and MHEALTH datasets. On the MM-Fit dataset, we additionally conducted experiments with pose, video, and multimodal inputs. The benefits of pre-training MuJo on FiMAD, especially when limited data is available for the target task, are analyzed. Furthermore, we analyze the zero-shot performance of our method and compare unimodal versus multimodal training on all four datasets.
@@ 2 Related Work
Unimodal and multimodal activity recognition.
Unimodal activity recognition has been extensively studied in HAR. Various approaches have been proposed to address this task, and they can be broadly categorized into two groups: hand-crafted feature-based and deep learning-based methods. Hand-crafted feature-based methods rely on extracting low-level features from the input data and designing a classifier to recognize activities. For instance, in accelerometer-based activity recognition Kwapisz et al. (2011), features such as mean, standard deviation, and energy of the acceleration signal are often used to capture the characteristics of different activities. Popular classifiers used in this approach include decision trees, support vector machines, and k-nearest neighbors. Deep learning-based methods, on the other hand, use neural networks to learn features and recognize activities from raw input data. Convolutional neural networks (CNNs) and recurrent neural networks (RNNs) are two popular types of neural networks used in this approach. CNNs are commonly used for image-based activity recognition Strömbäck et al. (2020), being combined with RNNs in some cases Ordóñez and Roggen (2016). Despite the success of unimodal activity recognition, it is often limited by the fact that a single modality of input data may only capture some of the necessary information about the activity, leading to lower recognition accuracy. Therefore, multimodal activity recognition, which combines information from multiple sources, has gained increasing attention in recent years.
Multimodal activity recognition, on the other hand, recognizes human activities using information from multiple modalities, such as IMU data, audio, and video Mekruksavanich and Jitpattanakul (2022); Ijaz et al. (2022). These features can then be combined using fusion methods such as early fusion or late fusion Duhme et al. (2022). Some studies have also focused on domain adaptation, where the goal is to recognize activities in a new environment using data from a different environment Plananamente et al. (2022). In such cases, transfer learning techniques can adapt the model to the new environment using a small amount of labeled data. Overall, multimodal activity recognition has shown promising results, with improved accuracy compared to unimodal approaches. However, challenges in data fusion, feature extraction, and model selection still need to be addressed to improve the performance of multimodal activity recognition systems Islam et al. (2022).
@@
Different modalities, e.g., natural language, visual inputs, or sensor signals, often complement a common concept. A key challenge of multimodal feature representation is exploring efficient methods for multimodal fusion with the given modalities to avoid missing modal features. We can differentiate the representation learning methods in two aspects: learning objectives and fusing the learned features in a joint space. In learning objectives, various approaches have been proposed to learn multimodal representations by formulating proxy tasks such as reconstruction, completion Seo et al. (2021); Sun et al. (2019), matching/alignment Li et al. (2020); Miech et al. (2020), and ordering tasks Li et al. (2020). Generating one feature vector with multimodal inputs for a common concept is a crucial problem of multimodal representation learning. Simple methods, such as concatenating learned uni-modal feature vectors, have been extensively explored to enhance the performance of downstream tasks Hu et al. (2018); Strömbäck et al. (2020). These methods involve combining feature vectors from a single modality, such as text or image, to create more informative representations that can be used for various tasks. More advanced methods have also been proposed, including a tensor fusion network Zadeh et al. (2017) and a transformer-based co-attention mechanism for efficient cross-modality learning Lu et al. (2019).
Proxy task were also developed for the field of Human Activity Recognition with weareable devices including IMUs, where unlabelled IMU data is used to pre-train better representations. Those approaches include Multitask Saeed et al. (2019), Auto-encoder Haresamudram et al. (2019) and SimClr Tang et al. (2021). Multi-task Saeed et al. (2019) proposes auxiliary tasks of binary classification for extracting useful features for the downstream task. A multi-task temporal convolutional neural network (CNN) is trained to classify whether the input signal was transformed or not for each of a set of transformations such as scaling, rotating the signal, etc. The authors utilize a large unlabelled accelerometer dataset for self-supervised learning and transfer learned knowledge to a small, labelled activity recognition dataset. Auto-encoder Haresamudram et al. (2019) is the application of auto-encoders where an encoder network reduces the dimensionality of the data while a decoder one brings the representation back to its original dimensions, with the whole system learning to reconstruct the original signal. SimClr Tang et al. (2021) learns general representations from a dataset where a model is trained to match different views instances created from the same signal by contrasting them. The positive pair is obtained by the transformation of the input window, while the negative pair is formed by the rest input windows of the same batch. Then, the contrastive loss pulls the positive pair together while pulling the negative pair far away from each other in the representation space.
@@
Contrastive learning is as an efficient self-supervised framework applied across multiple domains, which learns similar/dissimilar representations from data that are organized into similar and dissimilar pairs. Recently, Radford et al. presented a novel approach, known as CLIP Radford et al. (2021), to jointly represent images and text descriptions by leveraging contrastive loss with the similarity of image-text pairs. CLIP learns two feature spaces with two encoders, i.e., image and text encoders, then projects the learned features into a shared latent space with contrastive loss. Contrastive learning has been extended to other modalities as well. For example, video-text data has attracted interest in studies focused on the shared space of video-text features Xu et al. (2021); Xue et al. (2022) and text-to-video generation Singer et al. (2022). MotionCLIP Tevet et al. (2022) trained an encoder to find the appropriate embedding of an input sequence in CLIP space and a decoder to generate the most fitting motion for a given CLIP space. Following this, IMU2CLIP Moon et al. (2022) was proposed to learn a joint space that aligns IMU motion sensor recordings with video and text by projecting them into a shared representation space.
@@
Studies in action recognition Li et al. (2020); Wang et al. (2021); Sun et al. (2022); Morshed et al. (2023) have also significantly enhanced the field of fitness activity recognition, despite traditionally focusing on a broad spectrum of actions such as “cutting the potato” and “cleaning the drawer”. Certain investigations in action recognition also endeavor to acquire knowledge of the feature space through contrastive learning, as exemplified by initiatives like HERO Li et al. (2020) and ActionCLIP Wang et al. (2021).
@@
The achievements discussed in the aforementioned studies have shown that contrastive learning has a significant impact on extracting a joint space with similar and dissimilar pairs. Consequently, the contrastive learning methodology employed in our study aims to acquire knowledge of the feature space pre-training to fitness activities. Contrary to previous work, we extend this line of work on contrastive learning to a unique multimodal setting that utilizes more than two modalities, such as video, IMU sensors, poses, and text, within a multitask training manner in an end-to-end training paradigm similar to ImageBind Girdhar et al. (2023). Thus, our model and data provide a cross-modal classification opportunity.
3 Dataset Collection and Construction
Annotating video descriptions is time-consuming and expensive. Thus, we create our dataset from a comprehensive and varied collection of videos with temporally localized chapter information, including chapter titles and start times, specifically for fitness activities. YouTube provides chapters within a video, each with its preview, providing information and context for different parts of the video. Those chapters are manually added by the uploader or automatically generated. Each chapter is a continuous, non-overlapping segment that divides the video into a video clip and its corresponding textual description.
@@ 3 Dataset Collection and Construction
We manually collect 1388 fitness videos with pre-annotated video chapters. Additionally, we retrieve the captions automatically generated by YouTube. All our videos show a tutor who teaches how to do fitness exercises (e.g., “squat with dumbbells”, “sit-ups”, “side deep squats”, etc.). The trainers are primarily performing HIIT home workouts, focusing on bodyweight exercises with no or less equipment. The exercise videos provide instructions for the activities while showing the activities clearly. The annotation does not include predefined categories or labels. Each video owner annotates the activity segment of the videos with their personal preferences, such as “deep squats with dumbbells”, “high jump squats”, etc., instead of only “squats”. Based on the temporal boundaries of the chapters within each video, we split the videos in shorter activity clips along with their corresponding caption segments. Since workout exercises are typically short, activity videos that are too long often contain multiple exercises and, thus, incorrect chapter annotations. Hence, we only keep exercises shorter than 2 minutes and discard longer exercises. After that, the dataset consists of 10,695 (video, chapter, caption)-samples.
@@
We convert all 10,695 exercise videos to 720p resolution and 50 frames per second via FFmpegs adaptive overlapped block motion compensation with bidirectional motion estimation. To enhance the dataset with diverse modalities, we use IMUTube Kwon et al. (2021) to obtain pose and virtual sensor (accelerometer and gyroscope) features. As a result, our dataset includes (video, pose, sensor, text)-pairs for each input. For video and text we extracte multiple features. Video features include using HERO Li et al. (2020), which is utilizing either CLIP Radford et al. (2021) or S3D Xie et al. (2018) embeddings. For text, we apply several normalization strategies to the raw chapters such as replacing abbreviations within the text (e.g., changing “l” to “left” and “r” to “right”) as well as removing stop words, numbers, and a few special symbols from a predefined list. Further, we use GPT-3.5 Turbo to generate detailed descriptions of the exercises. We then extract multiple text features with various models, including OpenAI’s text-embedding-ada-002 embedding model (ADA), CLIP, BERT, as well as language features and motion tokens extracted with the Text2Motion Transformer Guo et al. (2022). We apply IMUTube to generate virtual gyroscope and accelerometer data for sensors placed on all 16 joint nodes. We also use the implementation of IMUTube to extract 3D poses from the videos, which utilizes OpenPose followed by VideoPose3D for 3D pose lifting. Since IMUTube fails to extract sensor data for 884 exercise videos, we exclude these sequences. We use the remaining 9811 exercises as our dataset and split it into 8,227 train and 1,584 validation activity clips. The split is based on the level of the 1,388 original fitness videos, to ensure video independence. We do not create a test set because the dataset is used solely for pre-training, not for testing. Since workout session videos typically start and end with explanations given by the trainer, we discard the initial and final 15%222This value has proven to be effective in experimental investigations. of each activity clip. Despite these filtering steps, we receive a large dataset totaling around 78 hours in length. Instances are generated using sliding windows of 100 frames in length (i.e., 2 seconds) with 50 frames overlap from the videos. We found this size suitable for fitness exercises as a 2 second window generally captures enough meaningful information regarding the subject’s movements. In summary, our dataset comprises a large collection of short fitness activity clips containing video features, poses, virtual IMU data, chapter labels, captions, and GPT-generated descriptions as well as several types of features extracted from these text modalities.